OpenStudy (anonymous):

m^2-25/2 m = 5/2solve by completing the square

OpenStudy (anonymous):

hello panther

OpenStudy (anonymous):

m^2 - 25/2 m = 5/2

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

m^2-25/2m-5/2=0 m^2-25/2m+625/16 -665/16=0 (m-25/4)^2=665/16 taking root... m-25/4=+-6.44 m=12.56 m=-0.32

OpenStudy (anonymous):

it says you can use radicals if needed and commas to separate answers